OpenStudy (anonymous):

Solve this system of equations. Write the answer as an ordered pair. -2x + 6y = 6 -7x + 8y = -5

OpenStudy (a1234):

Use elimination. -2x + 6y = 6 -7x + 8y = -5 -8x + 24y = 24 -21x + 24y = -15 Subtract. 13x = 39 x = 3 Substitute into one equation. -2x + 6y = 6 -2(3) + 6y = 6 -6 + 6y = 6 6y = 12 y = 2 (3, 2) is the solution. Check to see if that works.

OpenStudy (anonymous):

iight i will, thank you.

OpenStudy (a1234):

You're welcome.